# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v1

@@ Abstract
Reconstruction attacks and defenses are essential in understanding the data leakage problem in machine learning. However, prior work has centered around empirical observations of gradient inversion attacks, lacks theoretical groundings, and was unable to disentangle the usefulness of defending methods versus the computational limitation of attacking methods. In this work, we propose a strong reconstruction attack in the setting of federated learning. The attack reconstructs intermediate features and nicely integrates with and outperforms most of the previous methods. On this stronger attack, we thoroughly investigate both theoretically and empirically the effect of the most common defense methods. Our findings suggest that among various defense mechanisms, such as gradient clipping, dropout, additive noise, local aggregation, etc., gradient pruning emerges as the most effective strategy to defend against state-of-the-art attacks.
${}^{\dagger}$ Stanford University  ${}^{\ddagger}$ New York University
@@ 1 Introduction
Machine learning research has transformed the technical landscape across various domains but also raises privacy concerns potentially (Li et al., 2023; Papernot et al., 2016). Federated learning (Konečný et al., 2016; McMahan et al., 2017), a collaborative multi-site training framework, aims to uphold user privacy by keeping user data local while only exchanging model parameters and updates between a central server and edge users. However, recent studies on reconstruction attacks (Yin et al., 2021; Huang et al., 2021), also known as gradient inversion attacks, highlight privacy vulnerabilities even within federated learning. Attackers can eavesdrop on shared trained models and gradient information and reconstruct training data using them. Even worse, honest but curious servers can inadvertently expose training data by querying designed model parameters. Some defending methods are also proposed and analyzed.
@@
Prior data reconstruction attacks have centered around gradient inversion (Zhu et al., 2019), attempting to match gradients generated from dummy and real input variables. Some variants in this line incorporate prior input data knowledge as regularization in the objective function (Jeon et al., 2021; Yin et al., 2021). As illustrated in Fig. 1, the reconstruction attack can be viewed as an inverse problem; from the observations $G$ that is generated from the unknown data and a known function (model’s gradient), the adversary intends to reconstruct the input data, and a defending method is to prevent this from perturbing the observations. To evaluate the effectiveness of a defense is to evaluate how robust it is to the strongest attack within the threat model.
However, previous attacks lack strength and theoretical guarantees, making the evaluation of the defense’s strength untrustworthy. Empirical understanding of attacking and defending methods is affected by factors like optimization challenges of non-convex objectives. Under the heuristic attacking methods from prior work, it is hard to disentangle the following two possibilities: 1) the algorithmic or computational barrier of the attack, and 2) the success of the defense. In this work, we will investigate the problem by introducing stronger attacks, as well as bringing more theoretical perspectives.
A line of theoretical work analyzing privacy attacks is differential privacy (DP) (Dwork, 2006), which guarantees privacy in data re-identification or membership inference. However, there are scenarios when data identity information is not sensitive but data itself is, and DP is not suitable to interpret data reconstruction attacks Guo et al. (2022). As a concrete example, when the observation $G$ is the sum of the samples, it yields no DP guarantee but recovering data from the sum is impossible as long as there are multiple data points. Although there are previous works discussed the reconstruction attack under differential privacy setting (Guo et al., 2022; Stock et al., 2022), their setting is different from federated learning where all samples but one are known and the goal of the attacker is to reconstruct the last data point, which is unrealistic in the federate learning setting. To this end, with DP not applicable in this setting, there is limited theoretical work and it is urgent to call for more fundamental investigations on the data reconstruction problems.
Recent work by Wang et al. (2023) introduced a theoretically sound feature reconstruction attack, but lacked practicality as it focused on two-layer fully connected neural networks, while for deep networks it had to manipulate the bottom layers’ weights in a transparently detectable way. We enhance this attack by combining gradient and intermediate layer matching, where we utilize the methodology in (Wang et al., 2023) as a pre-processing step of intermediate layer reconstruction. In the setting of an honest-but-curious server, our stronger attack demonstrates robustness to various defenses through empirical validation. Moreover, the robustness can be theoretically analyzed, differing from previous attacks. Under the framework of this attack, we can conduct a more systematic and scientific evaluation of privacy defenses’ utilities. Our main contributions are the following:
  • We propose a new gradient inversion attack method. Leveraging the use of an untrained U-Net as an image prior and intermediate feature matching. The proposed method can recover data from gradients and be robust to a wide range of defense methods.
  • We establish theoretical analyses to thoroughly evaluate the effectiveness of a wide range of common defense strategies against feature recovery, providing a theoretical perspective in reconstruction attack and defense.
  • We further conduct extensive experiments on the common defense strategies with the proposed attack method, connecting theoretical principles with practice.
@@ 1.1 Related Work
In federated learning (Konečný et al., 2016; McMahan et al., 2017), a line of work in gradient inversion aims to reconstruct training data from gradients. An early work by Aono et al. (2017) theoretically showed that reconstruction is possible under a single neuron setting. Following this line, Zhu et al. (2019) proposed the gradient matching framework, where the training data and labels are recovered by minimizing the $l_{2}$ distance between gradients generated by dummy variables and real data. Instead of using $l_{2}$ distance, Wang et al. (2020) proposed a Gaussian kernel-based distance function, and Geiping et al. (2020) used cosine similarity.
@@
Apart from using different distance measures, recent works introduced image priors to gradient matching. Geiping et al. (2020) added a total variation regularization term since pixels of a real image are usually similar to their neighbor. Wei et al. (2020a) introduced a label-based regularization to fit the prediction results. Balunovic et al. (2021) formulated the gradient inversion attack in a Bayesian framework. Yin et al. (2021) used a batch-normalization regularization to fit the BN statistics and a group consistency regularization to indicate the spatial equivariance of CNN. Instead of using regularization terms, Jeon et al. (2021) represented dummy training data by a generative model to preserve the image prior. In this method, the latent variables and the parameters of the generative model are trained simultaneously.
@@
Some recent works investigated new data reconstruction methods outside the framework of gradient matching. Wang et al. (2023) theoretically showed that the training data can be reconstructed from gradients of networks with random weights using tensor decomposition methods, where optimization is not needed. Haim et al. (2022) proposed a reconstruction method using the stationary property for trained neural networks, induced by the implicit bias of training.
@@
Defense strategies in this setting, however, are relatively less studied. The original setting of federated learning proposed by McMahan et al. (2017) used local aggregation to update multiple steps instead of gradients directly. Bonawitz et al. (2016, 2017) proposed a secure aggregation that can collect the gradients from multiple clients before updating. Geyer et al. (2017); Wei et al. (2020b) introduced differential privacy into federated learning, perturbing the gradients. Aono et al. (2017) demonstrated an encryption framework in the setting of federated learning. Moreover, many tricks designed to improve the performance of neural networks are effective in defending against privacy attacks, such as Dropout (Hinton et al., 2012), gradient pruning (Sun et al., 2017), and MixUp (Zhang et al., 2017). The effect of defending strategies against gradient matching attacks are discussed in (Zhu et al., 2019; Geiping et al., 2020; Huang et al., 2021).
2 A Stronger Attack for Defense Methods Evaluation
Prior research has demonstrated the potential to retrieve input directly from gradient values, specifically for image classification endeavors, by presenting it as an optimization problem: when provided with a neural network characterized by parameters $\theta$ and its associated gradient $\nabla_{\theta}\mathcal{L}(x^{\ast},y^{\ast})$ computed with a private data batch $(x^{\ast},y^{\ast})\in\mathbb{R}^{b\times d}\times\mathbb{R}^{d}$, $b$ is the batch size, $d$, for example, is the image size. Attack methods tries to recover $x$, an estimation of $x^{\ast}$:
$\arg\min_{x}\mathcal{L}_{\rm grad}(x;\theta,\nabla_{\theta}\mathcal{L}_{\theta}(x^{*},y^{*}))+\alpha\mathcal{R}_{\rm aux}(x)$ (1)
The optimization objective is twofold: the initial term $\mathcal{L}_{\rm grad}(x;\theta,\nabla_{\theta}\mathcal{L}_{\theta}(x^{\ast},y^{\ast}))$ ensures that the gradient of the retrieved batch $x$ aligns with the given gradients. Meanwhile, the regularization term $\mathcal{R}_{\rm aux}(x)$ refines the recovered image using established image prior(s) such as total variation. This is because Eq. (1) is non-convex, and inverting gradients is often under-determined, i.e. there exists a pair of different data having the same gradient even when the learning model is large.
Meanwhile, a recent work by Wang et al. (2023) introduced a novel tensor-based approach for reconstructing input data from the gradients of two-layer fully connected neural networks. Distinct from traditional optimization-based gradient matching techniques, this tensor-based strategy has been shown, both theoretically and empirically, to offer superior performance. In the pursuit of building a more robust attack method that enhances defense mechanisms evaluation. we explore the integration of this tensor-based methodology with existing optimization-based gradient matching approaches. Specifically, the tensor-based technique is applied to reconstruct intermediate features before the fully connected layers, which commonly serve as classification heads in modern neural network architectures. High-quality reconstruction of these features, as evidenced by (Wang et al., 2023), provides additional information that can be utilized to further regularize the objective equation Eq. (1), thereby making gradient matching more computationally tractable. As we demonstrated in Figure 2 and Table 1, introducing feature matching to gradient matching results in a more potent attack methodology. This integrated approach not only strengthens the attack but also facilitates a more comprehensive theoretical examination of various defense strategies.
2.1 Our attack algorithm
In this section, we present a stronger attack algorithm designed to rigorously evaluate various defense mechanisms. The method follows (Geiping et al., 2020) to use the same selection of $\mathcal{L}_{\rm grad}$ and research (Yin et al., 2021) for regularization grounded in batch normalization. A key innovation of our method is the integration of additional regularizations, inspired by the recent contributions of (Wang et al., 2023). This novel incorporation seeks to strength the attack strategy, ensuring a more precise and efficient evaluation of defense methods.
Feature reconstruction with tensor-based method
To utilize the feature matching regularization, we have to obtain the intermediate feature as a preprocessing step. Since the federated learning model cannot directly access the intermediate feature, a feature reconstruction with small errors is needed. The setting of recovering the intermediate feature is the same as reconstructing input data from the gradient of a fully connected 2-layer neural network. Specifically, we denote the input dimension by $d$ and the number of classes by $K$. For a two layer neural network $f({\bm{x}};\Theta)=\frac{1}{m}A\sigma(W{\bm{x}})$ with $m$ hidden nodes, where $W=[{\bm{w}}_{1},\dots,{\bm{w}}_{m}]\in\mathbb{R}^{m\times d}$, $A=[{\bm{a}}_{1},\dots,{\bm{a}}_{K}]^{\top}\in\mathbb{R}^{K\times m}$ and $\sigma$ is point-wise, denote $z_{k}=a_{k}^{\top}\sigma(W{\bm{x}})$ and $p_{k}=\mathrm{Softmax}(z_{k})$. Then with cross-entropy loss, for input ${\bm{x}}$ we can observe the partial derivative
$\frac{\partial\ell}{\partial a_{kj}}=\frac{\partial\ell}{\partial z_{k}}\frac{\partial z_{k}}{\partial a_{kj}}=\frac{1}{m}(p_{k}-y_{k})\sigma({\bm{w}}_{j}^{\top}{\bm{x}})$
from gradient. We adopt the tensor-based reconstruction method proposed by Wang et al. (2023). We can define $g({\bm{w}}_{j})=\sum_{k=1}^{K}c_{k}\frac{\partial\ell}{\partial a_{kj}}$, where $c_{k}$’s are some constant with carefully chosen such that $r_{i}:=\sum_{k=1}^{K}c_{k}(p_{ki}-y_{ki})$ is a constant. Then compute $\sum_{j=1}^{m}g({\bm{w}}_{j})H_{p}({\bm{w}}_{j}),$ where $H_{p}$ is the $p$-th Hermite function, and conduct tensor decomposition. See details of the method in Appendix B.
Strategic weight initialization.
Challenges in order recovery of the features.

### v2

@@ Abstract
Reconstruction attacks and defenses are essential in understanding the data leakage problem in machine learning. However, prior work has centered around empirical observations of gradient inversion attacks, lacks theoretical justifications, and cannot disentangle the usefulness of defending methods from the computational limitation of attacking methods. In this work, we propose to view the problem as an inverse problem, enabling us to theoretically, quantitatively, and systematically evaluate the data reconstruction problem. On various defense methods, we derived the algorithmic upper bound and the matching (in feature dimension and model width) information-theoretical lower bound on the reconstruction error for two-layer neural networks. To complement the theoretical results and investigate the utility-privacy trade-off, we defined a natural evaluation metric of the defense methods with similar utility loss among the strongest attacks. We further propose a strong reconstruction attack that helps update some previous understanding of the strength of defense methods under our proposed evaluation metric.
Stanford University   New York University  §$\S$ Peking University
@@ 1 Introduction
Machine learning research has transformed the technical landscape across various domains but also raises privacy concerns potentially (Li et al., 2023; Papernot et al., 2016). Federated learning (Konečný et al., 2016; McMahan et al., 2017), a collaborative multi-site training framework, aims to uphold user privacy by keeping user data local while only exchanging model parameters and updates between a central server and edge users. However, recent studies on reconstruction attacks (Yin et al., 2021; Huang et al., 2021) highlight privacy vulnerabilities even within federated learning. Attackers can eavesdrop on shared trained models and gradient information and reconstruct training data using them. Even worse, honest but curious servers can inadvertently expose training data by querying designed model parameters. Some defending methods are also proposed and analyzed (Zhu et al., 2019; Wei et al., 2020b; Zhang et al., 2018).
However, previous research on reconstruction attacks and defenses centered around empirical studies and lacked theoretical guarantees. Moreover, those experiments over defenses usually focus only on one specific attack, making the evaluation of defenses’ strength untrustworthy. Empirical understanding of attacking and defending methods is affected by factors like optimization challenges of non-convex objectives. Under the heuristic attacking methods from prior work, it is hard to disentangle the following two possibilities: 1) the algorithmic or computational barrier of the specific attack and 2) the defense’s success. Thus, a systematical evaluation of defenses involving information theoretical analysis and eliminating the attack algorithm’s computational barrier is needed.
A line of theoretical work analyzing privacy attacks is differential privacy (DP) (Dwork, 2006), which guarantees privacy in data re-identification or membership inference. However, there are scenarios when data identity information is not sensitive, but data itself is, and DP is unsuitable for interpreting data reconstruction attacks (Guo et al., 2022; Hayes et al., 2024). Although some previous works discussed the reconstruction attack with the more tailored Renyi-DP (Guo et al., 2022; Stock et al., 2022), their setting is when all samples but one are known, and the attacker’s goal is to reconstruct the last data point, which doesn’t apply to the federated learning setting. We will discuss how (Renyi-)DP is too stringent for the data leakage problem under federated learning in Section 2. Some underlying structures of the observation can yield no (Renyi-)DP guarantee but prevent the attacker from reconstructing the data.
@@
In our work, we instead propose a theoretical framework that views the data reconstruction as an inverse problem (Fig. 1): from the gradient $G$ that is generated from the unknown data $S$ and a known function (training loss’s gradient), the adversary intends to reconstruct the input data, and a defender $D$ is to prevent this from perturbing the observations. Our theoretical analysis focuses on 2-layer fully connected networks and gives both upper and lower bounds on the error of this inverse problem under various defenses. The algorithmic upper bounds are based on a specific attack algorithm proposed by (Wang et al., 2023), and the lower bounds are derived from a statistical point of view. To generalize the analysis to real-world models and study the utility-privacy trade-offs, we empirically conduct a comprehensive evaluation of different defenses’ performances. Our comparisons are based on the strongest attack corresponding to each defense, so the result is general instead of especially effective for a specific attack.
Our paper is organized as follows. In Section 2, we introduce the exact setups of both theoretical and empirical analysis of defenses. We also explain why DP is not suitable for our setting. In Section 3, we summarize the defenses used in reconstruction attacks and analyze the error bounds theoretically for 2-layer networks. In Section 4.1, we introduce several previous attacking methods used in our empirical analysis and propose a strong attack based on an intermediate feature reconstruction method (Wang et al., 2023). In Section 5, we show the results of experiments and evaluate the defenses based on utility loss and best reconstruction error.
@@ 1.1 Related Work
In federated learning (Konečný et al., 2016; McMahan et al., 2017), a line of work in gradient inversion aims to reconstruct training data from gradients. Early work by Aono et al., 2017 theoretically showed that reconstruction is possible under a single neuron setting. Zhu et al., 2019 next proposed the gradient matching framework, where the training data and labels are recovered by minimizing the $l_{2}$ distance between gradients generated by dummy variables and real data. Instead of using $l_{2}$ distance, Wang et al., 2020 proposed a Gaussian kernel-based distance function, and Geiping et al., 2020 used cosine similarity.
@@
Besides using different distance measures, recent works introduced image priors to gradient matching. Geiping et al., 2020 added a total variation regularization term since pixels of a real image are usually similar to their neighbor. Wei et al., 2020b introduced a label-based regularization to fit the prediction results. Balunovic et al., 2021 formulated the gradient inversion attack in a Bayesian framework. Yin et al., 2021 used a batch-normalization regularization to fit the BN statistics and a group consistency regularization to indicate the spatial equivariance of CNN. Instead of using regularization terms, Jeon et al., 2021; Xue et al., 2023; Li et al., 2022 represented dummy training data by a generative model to preserve the image prior.
@@
Some recent works investigated new data reconstruction methods outside the framework of gradient matching. Wang et al., 2023 and Kariyappa et al., 2023 respectively use tensor-based method and independent component analysis to construct data from gradient information, where optimization is not needed. Haim et al., 2022 proposed a reconstruction method using the stationary property for trained neural networks induced by the implicit bias of training. Other works investigate partial data reconstructions with fishing parameters (Wen et al., 2022; Boenisch et al., 2023; Fowl et al., 2021).
@@
Defense strategies in this setting, however, are relatively less studied. The original setting of federated learning proposed by McMahan et al., 2017 used local aggregation to update multiple steps instead of gradients directly. Bonawitz et al., 2016, 2017 proposed a secure aggregation that can collect the gradients from multiple clients before updating. Geyer et al., 2017; Wei et al., 2020a introduced differential privacy into federated learning, perturbing the gradients. Aono et al., 2017 demonstrated an encryption framework in the setting of federated learning. Moreover, many tricks designed to improve the performance of neural networks are effective in defending against privacy attacks, such as Dropout (Hinton et al., 2012), gradient pruning (Sun et al., 2017), and MixUp (Zhang et al., 2017). The effect of defending strategies against gradient matching attacks are discussed in (Zhu et al., 2019; Geiping et al., 2020; Huang et al., 2021).
2 Preliminary
In reconstruction attacks, we observe the model iterations in training and recover data from the gradient $\nabla L(\Theta;S)$, where $L$ is the loss function, $\Theta$ is the model and $S$ is an unknown dataset. Then, it can be regarded as an inverse problem to identify unknown signal $S$ from observation $G=\nabla L(\Theta;S)$, where $\nabla L$ is the forward function. We denote the data space by $\mathcal{X}$ and the batch size of data by $B$. Then $S\subset\mathcal{X}^{B}$. Under the inverse problem framework, the effectiveness of attacks and defenses becomes the feasibility and hardness of recovering $S$. Formally, we define a minimax risk of reconstruction error:
$R_{L}=\left(\min_{\hat{S}=\hat{S}(G)}\max_{S\subset\mathcal{X}^{B}}\min_{\pi}\operatorname{\mathbb{E}}\left[d(S,\pi(\hat{S}))\right]\right)^{1/2}.$ (1)
Here $d(S,\pi(\hat{S}))=\frac{1}{B}\sum_{i=1}^{B}\|S_{i}-\hat{S}_{\pi(i)}\|^{2}$, where $\pi$ is a permutation of $[B]$. We will study in Section 3 the reconstruction error lower bounded in the setting of a two-layer network with random weights ***Throughout the paper, we use a unified scaling for data and network weights similar to Mean-field view (Mei et al., 2018).. For upper bounding reconstruction errors, we consider the error for specific algorithm $A$. We define the reconstruction error by
$R_{U}=\max_{S\subset\mathcal{X}^{B}}\min_{\pi}\left(d(S,\pi(\hat{S}))\right)^{1/2}.$ (2)
Note that in both lower and upper bounds, we consider the minimum over permutation. It is because batched gradient descent adds gradient together in an unordered manner, so only identifying the set of data points without their correspondence is important.
Our studies are meant to complement the esteemed DP (Dwork, 2006) to analyze the reconstruction as DP is too strong and sometimes unnecessary for our studied settings. On one hand, scenarios exist when reconstruction is impossible, even when DP does not holdFor instance, one can easily verify that with linear net and $l_{2}$ loss, the observation becomes $\sum_{i=1}^{B}x_{i}$, where no DP is satisfied but individual recovery of $x_{i}$ is impossible for $B>1$ unless specific prior knowledge is assumed.. In addition, the price of achieving DP in the data reconstruction setting is high. For DP-SGD (Abadi et al., 2016), the well-established Gaussian mechanism requires a large variance that will destroy the utility of gradient information:
Proposition 2.1 (Short version of Proposition A.1).
For a two-layer neural network with $m$ hidden nodes and random weights, we denote the gradient by $G$. Under mild assumptions, the randomized mechanism $\mathcal{M}=G+\mathcal{N}(0,\sigma^{2}I)$ is $(\epsilon,\delta)-$ DP for any $\epsilon,\delta>0$ if $\sigma^{2}=\Omega(\frac{m\log(1/\delta)}{\epsilon})$.
In this work, we will theoretically and empirically evaluate defenses against reconstruction. For theoretical analysis, we bound the reconstruction error under various defenses $D$ by estimating Eq. (1) and (2), where the observation is $D(G)$ instead of $G$. The theoretical bounds will connect the model architecture, defenses’ strength, and data dimension. However, only theoretical analysis lacks an extension to general networks and utility-privacy trade-off. Therefore we introduce a systematic way to evaluate defenses by empirical analysis further. Within a set of defenses $\mathcal{D}_{\mathcal{U}}$ of the same level of utility loss $\mathcal{U}$, we will measure the following criterion for each $D\in\mathcal{D}_{\mathcal{U}}$:
$\mathcal{S}_{D}:=\max_{A\in\mathcal{A}}d(S,A(D(G))),$ (3)
which evaluates their strength against the most effective attacker in $\mathcal{A}$, the set of considered reconstruction attacks. In experiments, we conduct various attacks under different defenses to estimate Eq. (3). Then, we compare the best reconstruction error with the same utility loss for each defense.
3 Theoretical Analysis
In theoretical analysis, we mainly focus on two-layer neural networks $f({\bm{x}};\Theta)=\sum_{j=1}^{m}a_{j}\sigma({\bm{w}}_{j}^{\top}{\bm{x}})$ with $\ell$ be the square loss. Here $a_{j}\sim\mathcal{N}(0,1/m^{2})$ and ${\bm{w}}_{j}\sim\mathcal{N}(0,I)$. For the upper bound, we examine a specific attack method proposed by (Wang et al., 2023). The attack method involves the computation of gradient-related tensor $\sum_{j=1}^{m}g({\bm{w}}_{j})H_{p}({\bm{w}}_{j})$, where $g({\bm{w}}_{j}):=\frac{\partial\ell}{\partial a_{j}}$ and $H_{p}$ is the $p$-th Hermite function. By Stein’s lemma (Stein, 1981; Mamis, 2022), it is approximate to a tensor product $\sum_{i=1}^{B}c_{i}{\bm{x}}_{i}^{\otimes p}$ when ${\bm{w}}_{j}$’s are random Gaussian and we can conduct tensor decomposition to recover data.
Theorem 3.1 (Informal, Theorem 5.1 in (Wang et al., 2023)).
For a 2-layer network under mild assumptions, the reconstruction error of the tensor-based attack has an upper bound $R_{U}\leq\tilde{O}(B\sqrt{\frac{d}{m}})$ with high probability.
For the lower bound, to avoid pure combinatorial analysis, we formulate data reconstruction as a statistical estimation problem: we observe the noisy gradient $\nabla_{\Theta}L(\Theta;S)+\epsilon$ and identify the input data $S$. Here $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ is a Gaussian random vector. We consider the minimax lower bound in Eq. (1), where the expectation is over the random Gaussian noise and the high probability in the following proposition is derived with random weights.
Theorem 3.2 (Short version for Theorem D.7).
For a 2-layer network with noisy gradient, the statistical minimax risk has an upper bound $R_{L}\geq\tilde{\Omega}(\sigma\sqrt{\frac{d}{m}})$ with high probability.
In the following, we analyze the upper and lower bound for the reconstruction under different defenses. The bound is shown in Table 1. Our analysis provides a tight and quantitative analysis of how each defense’s defending strength (such as pruning ratio or noise variance) affects the data reconstruction error (which is inversely proportional to the degree of privacy). In the remaining part of this section, we individually investigate each defense method.
No defense Local aggregation Gradient noise Gradient clipping DP-SGD Dropout Gradient pruning
Upper bound $B\sqrt{\frac{d}{m}}$ $B\sqrt{\frac{d}{m}}$ $(B+\sigma_{0})\sqrt{\frac{d}{m}}$ $B\sqrt{\frac{d}{m}}$ $(B+\sigma_{0}\max\{1,\frac{\|G\|}{C}\})\sqrt{\frac{d}{m}}$ $B\sqrt{\frac{d}{(1-p)m}}$ Unknown
Lower bound $\sigma\sqrt{\frac{d}{m}}$ $\sigma\sqrt{\frac{d}{m}}$ $\sigma\sqrt{\frac{d}{m}}$ $\sigma\max\{1,\frac{\|G\|}{C}\}\sqrt{\frac{d}{m}}$ $\sigma\max\{1,\frac{\|G\|}{C}\}\sqrt{\frac{d}{m}}$ $\sigma\sqrt{\frac{d}{(1-p)m}}$ $\sigma\sqrt{\frac{d}{(1-\hat{p})m}}$
Table 1: The algorithmic upper bound and information-theoretic lower bound of the reconstruction error against different defenses. Some parameters are defined later in the subsection of each defense.
3.1 Local aggregation
In a realistic setting of federated learning, the observed local update may not be a single gradient. Instead, it can be the result of multiple steps of gradient descent (McMahan et al., 2017). In our analysis, we mainly consider the most simple cases, where local devices train two steps and the observation is $\Theta^{(2)}-\Theta^{(0)}$. Since one step of training does not change the parameters too much, we can approximately assume that the two steps share a same set of parameters. Then the analysis is similar to that without defense.
Proposition 3.1 (Short version for Proposition C.1).
For a 2-layer network with mild assumptions, the reconstruction error of tensor based attack under local aggregation defense for 2 steps has a upper bound $R_{U}\leq\tilde{O}(B\sqrt{\frac{d}{m}})$ with high probability.
Proposition 3.2 (Short version for Proposition E.2).
For a 2-layer network with noisy gradient and local aggregation defense with two steps, the statistical minimax risk has a lower bound $R_{L}\geq\tilde{\Omega}(\sigma\sqrt{\frac{d}{m}})$ with high probability.
3.2 Differential privacy stochastic gradient descent
Differential privacy (Dwork, 2006) algorithms ensure the privacy of data by adding noise to gradient descent steps (Shokri and Shmatikov, 2015; Abadi et al., 2016; Song et al., 2013). In this way, it hides the exact local gradient from the global model. In differential private federated learning (Geyer et al., 2017; Wei et al., 2020a), each client clips the gradient and introduces a random Gaussian noise before updates to the cluster. In the setting of differential privacy stochastic gradient descent (DP-SGD), the update of $a_{j}$ is
$\tilde{G}_{j}=G_{j}/\max\left\{1,\frac{\|G\|}{C}\right\}+\epsilon_{0},\text{ where }\epsilon_{0}\sim\mathcal{N}(0,\sigma_{0}^{2}I).$
Note that under this setting, the information contained in the gradient is disrupted by gradient clipping and random noise. For the upper bound, we first analyze the effect of the two strategies separately and then combine them to bound the error of DP-SGD. If we conduct the tensor-based attack on clipped gradient $\tilde{G}$, there is no defensive effect no matter the clipping threshold and the reconstruction error bound $\tilde{O}(B\sqrt{\frac{d}{m}})$ is the same as that with no defense (even for the hidden terms). For noisy gradient $G+\epsilon_{0}$, we have the error bound $\tilde{O}((B+\sigma_{0})\sqrt{\frac{d}{m}})$, which is the same order to original case if $\sigma_{0}\leq O(\frac{B}{m})$.
Proposition 3.3 (Short version for Proposition C.13).
For a 2-layer network with mild assumptions, the reconstruction error of tensor-based attack under defense DP-SGD with clipping threshold $S$ and Gaussian noise with variance $\sigma_{0}^{2}$ has a upper bound $R_{U}\leq\tilde{O}((B+\sigma_{0}\max\{1,\frac{\|G\|}{C}\})\sqrt{\frac{d}{m}})$ with high probability.
In the analysis of statistical lower bound, there is already a Gaussian noise in the observed gradient so we only need to analyze the effect of gradient clipping. Similarly, the clipping changes the scaling of the random noise and the bound changes to $\tilde{\Omega}(\sigma\max\{1,\frac{\|G\|}{C}\}\sqrt{\frac{d}{m}})$. Therefore, gradient noise will be effective if the scaling of noise is large. However, this hurts the utility a lot.
3.3 Secure aggregation
In federated learning, local devices update gradients to the server individually. Secure aggregation (Bonawitz et al., 2016, 2017) is a method that clients $1,\cdots l$ can aggregate their gradients before the cluster, and the cluster can only access to aggregated gradient: $\tilde{G}=\frac{1}{B}\sum_{l=1}^{L}G_{l}B_{l},$ where $G_{l}$ and $B_{l}$ is the gradient update and batch size for $l$-th user respectively and $B=\sum_{l=1}^{L}B_{l}$. In this way, the global server is blocked from knowing each gradient, which prevents privacy leakage. Even though the server can reconstruct the training data, it cannot identify which user has the data. If we only reconstruct data without identifying its source, the aggregated gradient is equivalent to a gradient step with a large batch size $B$, and both bounds remain unchanged.
3.4 Dropout
Another method to improve the defending effect is dropout (Hinton et al., 2012; Srivastava et al., 2014), a mechanism designed to prevent overfitting. It has been empirically discovered in (Zhu et al., 2019; Huang et al., 2021) that it can defend against reconstruction attack. It randomly drops nodes in a network with a certain probability of $1-p$. In this way, some of the entries in the gradient turn to zero, introducing randomness to the gradient. Thus, it is natural to prevent data leakage with dropout. A two-layer fully connected network with a dropout layer can be seen as a model with an effective width equal to the number of nodes $m^{\prime}\approx(1-p)m$ that has not been dropped. Then $m$ changes into $(1-p)m$ for both upper and lower bounds.
3.5 Gradient pruning
Gradient pruning (Sun et al., 2017; Ye et al., 2020) is a method that accelerates the computation in training by setting the small entries in gradient to zero. Thus, the gradient will be sparse, which is similar to dropout. However, the key difference is that the remaining entries of the gradient in dropout are chosen randomly yet gradient pruning drops small entries. With dropping nodes by the pruning rules, the distribution of effective weights is not Gaussian, which violates a key assumption in upper bound analysis. Therefore, there is no evidence showing that the tensor-based method can reconstruct input data with small errors under gradient pruning defense. In contrary, we can prove a lower bound with the same order when there is no defense.
Proposition 3.4 (Short version for Proposition E.3).
[…]
No upper bound derived indicated gradient pruning is more effective for the tensor-based data reconstruction method. The lower bound also demonstrates its potential for better privacy-utility trade-off, at least compared to dropout. Note that gradient pruning only mildly hurts utility by pruning out the least important coordinates; it does not necessarily hurt reconstruction error in the same way.
4 Empirical Analysis
4.1 Attacks for Defense Methods Evaluation
To conduct a fair comparison between different defense $D$, we evaluate their strength in response to the strongest attack: $\mathcal{S}_{D}:=\max_{A\in\mathcal{A}}d(S,A(D(G)))$, where $\mathcal{A}$ is the set of attacking methods considered in this evaluation. The ideal case is for $\mathcal{A}$ to contain the strongest possible attack; to complement existent attacks, we propose a new attack demonstrated strong for various defenses. We also include a range of attacks that are either related to our proposed attack method or widely acknowledged to be strong and robust to different defenses:
  • GradientInversion (Geiping et al., 2020). We consider the attack proposed by (Geiping et al., 2020), which achieves good recovery results using gradient inversion with known BatchNorm statistics, as knowing such statistics often results in better recovery results and thus stronger attack (Huang et al., 2021).
  • Our proposed attack. Our proposed attack is an improved version of GradientInversion attack. There are two modules we introduced to the original GradientInversion framework. a) a randomly initialized image prior network is introduced to generate the images, and b) the latent features of the fully connected layer are recovered first based on (Wang et al., 2023), and then inverting these features using feature inversion to recover the input images. More details will be introduced in Section 4.2.
  • CPA (Kariyappa et al., 2023). We consider this attack method due to its similarity to our method: GradientInversion and feature inversion are both adopted to recover the input images. In contrast to our proposed attack, CPA forms the feature recovery problem as a blind source separation problem and recovers features by finding the unmixing matrix.
  • Robbing The Fed (Fowl et al., 2021). As both the proposed method and CPA require minimal malicious modifications of the shared model, we are also interested in attacks on explicitly malicious servers.
We apply various defense techniques to the above attack algorithms, and for each defense, we consider the best attack performance across different attack algorithms following Eq. 3.
4.2 Description of the proposed attack algorithm
Notice that the tensor-based reconstruction attack is provably strong, with a matching lower bound in the setting of two-layer neural networks. However, the original design in (Wang et al., 2023) was impractical, and we propose an attack that incorporates the strength of the tensor-based method and the generality of gradient inversion attacks for a comprehensive evaluation of different defenses. On top of the gradient matching term $\mathcal{L}_{\rm grad}$ in (Geiping et al., 2020) and the prior knowledge utilized in (Yin et al., 2021) grounded in batch normalization, the key innovation of our method is an additional regularization that integrates feature reconstruction to gradient inversion. However, the integration is challenging due to different parameter requirements for gradient inversion attacks and tensor-based feature reconstruction.